OpenStudy (anonymous):

a circle has a circumference of 18 pi cm. find the radius?

jimthompson5910 (jim_thompson5910):

C = 2*pi*r 18*pi = 2*pi*r 18*pi/pi = 2*pi*r/pi 18 = 2*r r = ??

OpenStudy (anonymous):

9?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you ! C:

jimthompson5910 (jim_thompson5910):

yw